What is the HS- ion concentration of a 0.1 F H2S solution which is also 0.1 F in present to HC2H3O2?



1
Expert's answer
2021-04-01T02:44:18-0400

ionization constant of acetic acid "= 1.8\u00d710^{-5}"


the first ionization constant of H2S (about "1.26\u00d710^{-7}," a datum not present in the problem).


For a 0.1 F aqueous solution of acetic acid only, the concentration of H3O+ ions is readily found to be about "1.34\u00d710^{-3} mol\/L."


Then you write the 1st ionization constant of H2S as "1.26×10^{-7 }= (1.34×10^{-3}+ x)× \frac{x}{(0.1 - x)} ,"


where x is the molar concentration of HS^- ions.


This is a 2nd degree equation in x. It can be reasonably assumed that x is very small with respect to "1.34\u00d710^{-3}" , and, a fortiori, with respect to 0.1 , so that, as a first approximation, you can write "1.26×10^{-7 }= (1.34×10^{-3} )×\frac{x}{(0.1)}" ,


which yields "x = 9.4\u00d710^{-6} mol\/L" .


By employing this first-approximation value of x, you rewrite "1.26×10^{-7} = (1.34×10^{-3} +( 9.4×10^-6)×\frac{x}{(0.1) }"


and find "x = 9.3\u00d710^{-6} mol\/L," a value that needs no further to be improved, on the basis of the significant figures present in the data of the problem. As a conclusion, the equilibrium concentration of HS^- ions should be about "9.3\u00d710^{-6} mol\/L."